OpenStudy (anonymous):

Find an equation of the line satisfying the following conditions. If possible, write the equation in slope-intercept form. Through the point (-6,-7) and parallel to 4x-3y=-36

OpenStudy (lalaly):

parallel means i has the same slope \[y=\frac{4x}{3}+12\] so the slope= \[\frac{4}{3}\] \[y-(-7)=\frac{4}{3}(x-(-6))\]\[y+7=\frac{4}{3}x+8\]\[y=\frac{4}{3}x+1\]